OpenStudy (anonymous):

The numbers 20, 21, and c are a Pythagorean triple. What is the value of c? A. 41 B. 29 C. 25 D. 22

OpenStudy (noseboy908):

When finding missing values here, the specified letter is crucial. Thinking of a right triangle, what are the two sides often called that border the right angle? What letters are used to describe them frequently?

OpenStudy (anonymous):

See I suck at the pathorgeon theory and I can't solve it... I dont understand

OpenStudy (praxer):

20^2+21^2=841=29^2 hence the side c is 29

OpenStudy (noseboy908):

Not a problem. Ignoring Praxor's blatant answer, the sides I've mentioned are amost always represented by the variables a and b. "c" is almost always the longest side. Now, knowing that, can you state the Pythagoras theorm please?

OpenStudy (anonymous):

Was he correct? @noseboy908

OpenStudy (anonymous):

a^+b^=c^

OpenStudy (praxer):

I am correct a^2+b^2=c^2 is the pythagorean theorem. http://en.wikipedia.org/wiki/Pythagorean_theorem

OpenStudy (anonymous):

Okay Thank you!

OpenStudy (praxer):

Welcome... :)

OpenStudy (noseboy908):

He is indeed correct.

OpenStudy (anonymous):

Evaluate the expression by using the given values of the variables. y ÷ 3 • 2 – 8 ÷ x + 5 where x = 4 and y = 12 A. 5 B. 9 C. 11 D. 21

OpenStudy (noseboy908):

Alright. To begin, you need to substitute your given x and y values in for x and y in your equation. Please do that, but don't solve anything yet. Just restate the equation with the new numbers in place of the variables please.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x would be 4 and y is 12

OpenStudy (noseboy908):

Yep. So fill those in.

OpenStudy (anonymous):

.... how would that be?

OpenStudy (noseboy908):

What I mean is to, in the place of y in the original equation, place what you're told y is equal to, 12. And do the same for x.

OpenStudy (anonymous):

oh oh oh like this 12 ÷ 3 • 2 – 8 ÷ 4+ 5

OpenStudy (anonymous):

and then i dont know after that

OpenStudy (anonymous):

the answer is 11

OpenStudy (noseboy908):

The most important step now is to use the order of operations correctly. And yes, you're correct.

OpenStudy (anonymous):

Thanks i have a few more

OpenStudy (noseboy908):

Ask away. I have a little time left.

OpenStudy (anonymous):



OpenStudy (anonymous):

A. –15 B. –10 C. 3 D. 8

OpenStudy (noseboy908):

Alright. To begin, our overall goal is to isolate the z. So, how do we move the two to the opposite side?

OpenStudy (anonymous):

multiply?

OpenStudy (noseboy908):

Yep. What do you get when you do?

OpenStudy (anonymous):

18

OpenStudy (noseboy908):

No, right after you multiply the two, what does the equality become?

OpenStudy (anonymous):

umm.. Could you show me?